# ProSGNeRF: Progressive Dynamic Neural Scene Graph with Frequency Modulated Auto-Encoder in Urban Scenes

### v2

@@ Abstract
Implicit neural representation has demonstrated promising results in view synthesis for large and complex scenes. However, existing approaches either fail to capture the fast-moving objects or need to build the scene graph without camera ego-motions, leading to low-quality synthesized views of the scene. We aim to jointly solve the view synthesis problem of large-scale urban scenes and fast-moving vehicles, which is more practical and challenging. To this end, we first leverage a graph structure to learn the local scene representations of dynamic objects and the background. Then, we design a progressive scheme that dynamically allocates a new local scene graph trained with frames within a temporal window, allowing us to scale up the representation to an arbitrarily large scene. Besides, the training views of urban scenes are relatively sparse, which leads to a significant decline in reconstruction accuracy for dynamic objects. Therefore, we design a frequency auto-encoder network to encode the latent code and regularize the frequency range of objects, which can enhance the representation of dynamic objects and address the issue of sparse image inputs. Additionally, we employ lidar point projection to maintain geometry consistency in large-scale urban scenes. Experimental results demonstrate that our method achieves state-of-the-art view synthesis accuracy, object manipulation, and scene roaming ability. The code will be open-sourced upon paper acceptance.
@@ 1 Introduction
Scene reconstruction and view synthesis on urban scenes have significantly transformed our lives, particularly in navigation and autonomous driving. NeRF [14] is the most impressive method using implicit neural representation and has shown promising results for view synthesis on static and object-centric scenes. Some recent works improve the original NeRF and achieve large-scale urban scene reconstruction. Block-nerf [19] decomposes the scene into multiple blocks and enables per-block update for virtual drive-through reconstruction. Mega-nerf [20] uses expanded octree architecture for large-scale photo-realistic fly-through scenarios. Bungeenerf [24] design a growing model with residual block structure for city-level scene reconstruction.
Although existing methods have made improvements, they are still insufficient for real-world urban scenes. We summarize several core challenges for urban scene reconstruction: a) multi-dynamic objects and large-scale ego-motions. Existing algorithms have limitations in handling scenes with both complex object motions and camera movements. Their representation capabilities are also limited with their single, global scene representation for the entire scenario. b) sparse view input for dynamic objects. In real-world urban scenes, we typically have only a few viewpoints for dynamic objects, which leads to poor reconstruction performance. c) lack of annotations and labeling It is also challenging to obtain annotations and labels for real-world scenes.
Figure 1: The isometric view of the proposed method, ProSGNeRF. We employed a 2D segmentation network, SAM, to preprocess the training data and generate accurate masks for dynamic objects. We propose a progressive neural scene graph approach that dynamically allocates local scene graph (box). The entire scene is decomposed into three parts: background, dynamic objects, and far-field. We design separate networks for background and objects and introduce a far-field loss for regularization. Nodes $l$ represent individual dynamic objects. $F_{bkg}$ models the static background scene and $F_{obj}$ models movable foreground objects in local object-centric coordinate frames.
@@
To this end, we propose ProSGNeRF for large-scale dynamic urban scene reconstruction and novel view synthesis. For the first challenge, we utilize a neural scene graph for learning the local scene representation with decoupled object transformations and background representation. The neural scene graph incorporates a progressive scheme that employs a temporal window strategy to effectively address scalability, dividing the entire scene into multi-local scenes and dynamically initiating and maintaining local scene graphs. For the second challenge, we design a frequency auto-encoder network to encode shape and appearance features for dynamic objects and regularize the frequency range of dynamic object information, which is efficient for sparse view input and enhances the representation of dynamic objects. For the third challenge, the Segment Anything Model (SAM) has demonstrated its remarkable ability to revolutionize 2D segmentation. We leverage the promptable segmentation method to predict masks for dynamic objects and far-field scenarios, which can reduce the need for human annotation. We also use lidar projection supervision for geometry consistency. To validate our proposed framework, we evaluate the effectiveness of the method on different outdoor urban datasets, such as KITTI and VKITTI datasets [8, 3], and do exhaustive evaluations and ablation experiments on these datasets. Our system demonstrates superior performance compared to concurrent methods [16, 21] that scene reconstruction on urban scenes. In summary, our contributions are shown as follows:
@@
  • We propose a novel editable scene reconstruction method, ProSGNeRF, which decomposes dynamic, background, and far-field into a progressively neural scene graph. Our method also supports object manipulation and scene roaming.
@@
  • The proposed ProSGNeRF method processes the large-scale scenes by dynamically instantiating local neural scene graphs, enabling scalability to handle arbitrary scale scenes.
@@
  • Due to the sparse observation of objects in the urban scene, we design a frequency auto-encoder architecture network to efficiently encode shape and appearance features for the objects of interest and perform frequency modulation for sparse view inputs.
@@ 2 Related work
Recently, with the introduction and advancement of Neural Radiance Fields [14], a broad array of subsequent research has been inspired, building upon the original approach. We review the related work in the areas of implicit scene representations, dynamics, and large-scale scene representation.
@@
Implicit Scene Representation  NeRF [14] is the most impressive work in view synthesis and 3D reconstruction. It proposes implicit representations to model scenes as functions, encoding the scene using a multi-layer perceptron (MLP). NeRF++ [26], mip-nerf [1], and mip-NeRF 360 [2] further improve the representation ability for unbounded environments. Those methods use the non-linear scene parameterization method to model unbounded scenes. NeRF-W [13] incorporate per-image embeddings to model appearance variation of the outdoor scenes.
@@
Dynamics  There are some methods that focus on dynamic object reconstruction, such as Neural 3D Video Synthesis [10] and Space-time Neural Irradiance Fields [23]. These approaches encode time information into scene representation. [11] and [7] incorporate 2D optical flow with warping-based regularization losses to enforce multi-view consistency and dynamic object perception. Nerfies [17] and Hypernerf [18] design a per-frame deformation field for dynamic humans. These methods are limited to single-object scenes, whereas our method scales beyond synthetic and indoor scenes to urban scenes .
@@
Large-scale Scene Representation For large-scale scene representation, [24] proposes Bungeenerf with a progressively growing model for city-level rendering. Block-NeRF [19] and Mega-NeRF [20] decompose the scene into multiple local radiance fields for large-scale scenes. However, these methods can only deal with static environments. Dynamic objects are prevalent in real-world urban road scenes. Panoptic-NeRF [6] and Panoptic Neural Fields [9] produce panoptic segmentation jointly with scene representations, achieving urban scene reconstruction.
@@
NSG [16], SUDS [21], and Mars [22] are the most relevant works to ours. Both NSG and Mars adopts neural scene graphs and dynamic object bounding boxes to decouple the scene but cannot handle ego-motion. Besides, NSG needs over 1TB of memory to represent a 30-second video. SUDS [21] uses a large model (DINO [4]) to distill scene features and obtain semantic information and scene flow. Their methods can handle camera ego-motion and static objects but struggle with high-dynamic objects. In contrast, our approach effectively addresses large-scale camera ego-motion and dynamic objects, enabling urban scene editing useful for constructing simulators for autonomous driving.
@@ 3 Method
The pipeline of our system is shown in Fig. 1. We use a set of sequential RGB images $\{I_{i}\}_{i=1}^{M}$ with known pose $\{R|t\}$ as our input. We construct the progressive neural scene graph with the instance mask from SAM and 3D bounding box. The scene is decomposed into multi-local scene graphs with three branches: background, dynamic objects, and far-field. In the dynamic objects branch, we design a frequency auto-encoder to learn a normalized, object-centric representation whose embedding describes and disentangles shape, appearance, and pose. The decoder translates the features into occupancy and color information for given 3D points and viewing directions in object space. We also design lidar points projection geometry consistency in large-scale urban scenes.
[…]
@@ 3.1 Progressive Neural Scene Graph
To this end, we design a progressive neural scene graph to represent city-scale scenes with multi-dynamic objects. In Figure 2, the scene graph $\mathrm{SG}$ is depicted, encompassing cameras, static nodes, and a collection of dynamic nodes. These dynamic nodes capture the evolving aspects of the scene, encapsulating the object’s appearance, shape, and various classes.
The entire scene is decomposed into several local scene representations:
$\{I_{i},D_{i}\}_{i=1}^{M}\mapsto\{\mathrm{SG}^{1}_{\theta_{1}},\mathrm{SG}^{2}_{\theta_{2}},\dots,\mathrm{SG}^{n}_{\theta_{n}}\}\mapsto\{c,\sigma\}$
@@
where the root node of scene graph denoted as $\mathcal{W}$, $C$ is a leaf node representing the camera and its intrinsic $K\in\mathbb{R}^{3\times 3}$. $L$ denotes the leaf node of scene graph(e.g. objects and background). $F$ represents the neural networks for both static and dynamic nodes: $F=F_{\theta_{bckg}}\cup\left\{F_{\theta_{1}},F_{\theta_{2}},\dots,F_{\theta_{c}}\right\}$. $E$ are edges that represent affine transformations from $i$ to $i^{\prime}$:
$E$ =$=\left\{e_{i,i^{\prime}}\in\left\{[],\boldsymbol{T}_{i\to i^{\prime}}\right\}\right\},$ (2)
@@
$\boldsymbol{R}_{i\to i^{\prime}}$ $\in\mathbb{R}^{3\times 3},\quad\boldsymbol{t}_{i\to i^{\prime}}\in\mathbb{R}^{3}\text{. }$
@@
In our method, we dynamically create local scene representation. Whenever the estimated camera pose trajectory reaches the bound of the current scene graph, we initialize the local representation with a local subset of frames, and then we progressively introduce subsequent frames to the optimization. Each subset contains some overlapping images, essential for achieving consistent reconstructions in the global scene. We further use inverse distance weight (IDW) fusion for all overlapping scene representations. Whenever the camera pose leaves the bound of the current scene representation, we stop optimizing previous ones (freeze the network parameters). At this point, we can reduce memory requirements and clear any supervising frames that are not needed anymore.

### v3

# ProSGNeRF: Progressive Dynamic Neural Scene Graph with Frequency Modulated Foundation Model in Urban Scenes

@@ Abstract
Implicit neural representation has demonstrated promising results in 3D reconstruction on various scenes. However, existing approaches either struggle to model fast-moving objects or are incapable of handling large-scale camera ego-motions in urban environments. This leads to low-quality synthesized views of the large-scale urban scenes. In this paper, we aim to jointly solve the problems caused by large-scale scenes and fast-moving vehicles, which are more practical and challenging. To this end, we propose a progressive scene graph network architecture to learn the local scene representations of dynamic objects and global urban scenes. The progressive learning architecture dynamically allocates a new local scene graph trained on frames within a temporal window, with the window size automatically determined, allowing us to scale up the representation to arbitrarily large scenes. Besides, according to our observations, the training views of dynamic objects are relatively sparse according to rapid movements, which leads to a significant decline in reconstruction accuracy for dynamic objects. Therefore, we utilize a foundation model network to encode the latent code. Specifically, we leverage the generalization capability of the visual foundation model DINOv2 to extract appearance and shape codes, and train the network on a large-scale urban scene object dataset to enhance its prior modeling ability for handling sparse-view dynamic inputs. In parallel, we introduce a frequency-modulated module that regularizes the frequency spectrum of objects, thereby addressing the challenge of modeling sparse image inputs from a frequency-domain perspective. Experimental results demonstrate that our method achieves state-of-the-art view synthesis accuracy, object manipulation, and scene roaming ability in various scenes. The code will be open-sourced on The code will be open-sourced on https://github.com/dtc111111/prosgnerf.
keywords:
3D Scene Reconstruction, Foundation Model,Neural Scene Graph, Large-scale Urban Scenes
Figure 1: Urban scene reconstruction and editing with ProSGNeRF. We show our view synthesis in different time steps (65,262) and scene decomposition results. Our approach significantly improves the view synthesis performance in real-world urban scenes containing multiple dynamic objects and large-scale camera ego-motion. We highlight and enlarge objects in the first column of images, providing the corresponding object PSNR in the top left corner. Scene PSNR is provided in the second column.
@@ 1 Introduction
Scene reconstruction and view synthesis on urban scenes have significantly changed our lives, particularly in autonomous driving [deng2025gaussiandwm3dgaussiandriving], city simulation, and virtual reality (VR) [fmgs] and Augmented Reality (AR) [cen2025segment]. The development of autonomous driving requires substantial and accurate testing to ensure safe behavior when deployed in the real world. View synthesis and controllable simulation for driving scenes also enable the generation of corner cases, and safety-critical situations aid in validating and enhancing the safety of autonomous driving systems at a lower cost.
NeRF [NeRF] is the most impressive method using implicit neural representation and has shown promising results for view synthesis on static and object-centric scenes. Some recent works improve the original NeRF and achieve large-scale urban scene reconstruction. Block-nerf [blocknerf] pre-divides the scene into multiple blocks with per-block update for virtual drive-through reconstruction. Mega-nerf [meganerf] uses expanded octree architecture for large-scale photo-realistic fly-through scenarios. Bungeenerf [bungeenerf] design a growing model with residual block structure for city-level scene reconstruction. [Localrf, mneslam] focus on incremental scene representation, where multiple MLP modules are employed to model the entire scene, and the representations are fused through an online distillation strategy. NSG [nsg] introduces a scene graph architecture to represent the scene, while SUDS [suds] factorizes the urban scene into three separate hash table data structures to efficiently encode static, dynamic, and far-field radiance fields. Mars [mars] proposes a modular framework with instance-aware mask and separately with independent networks so that the static model the dynamic properties of instances and background. EmerNeRF [emernerf] represents the scenes into static and dynamic fields. It also uses the flow field to aggregate multi-frame features, improving the rendering precision of dynamic objects. Although existing methods have made improvements, they are still insufficient for real-world urban scenes. We summarize several core challenges for urban scene reconstruction: a) multiple dynamic objects. Existing algorithms have limitations in handling scenes with complex multiple-object motions. b) insufficient scene representation capability. Their representation capabilities are also limited with their single, global scene representation for the entire scenario. c) sparse view input for dynamic objects. In practical autonomous driving scenarios, observations of the same vehicle are limited to only a few frames (e.g. four or five images for a car), which leads to poor reconstruction performance for dynamic objects.
@@
To this end, we propose ProSGNeRF for large-scale dynamic urban scene reconstruction and novel view synthesis. First, we propose a progressive neural scene graph architecture for learning the local scene representation with decoupled object transformations, background and far-field representation. The neural scene graph incorporates a progressive scheme that employs a temporal window strategy to effectively address scalability, dividing the entire scene into multi-local scenes and dynamically initiating and maintaining local scene graphs. In each local scene graph, we incorporate multiresolution hash encoding into the background representation to better model low-level architectural details. Second, we design a frequency foundation model network to encode shape and appearance features for dynamic objects and regularize the frequency range of dynamic object information, which is efficient for sparse view input and enhances the representation of dynamic objects. We ultilize the visual foundation model DINO V2 [dino] to extract appearance and shape codes for dynamic objects with the prior information to enhance the reconstruction accuracy of dynamic objects. Third, the Segment Anything Model [sam] has demonstrated its remarkable ability to revolutionize 2D segmentation. We leverage the promptable segmentation method to predict masks for dynamic objects and far-field scenarios, which can reduce the need for human annotation. We also use lidar projection supervision for geometry consistency. To validate our proposed framework, we evaluate the effectiveness of the method on different outdoor urban datasets, such as KITTI [kitti], VKITTI [vkitti], Waymo [waymo], nuScenes [nuScenes] datasets, and do exhaustive evaluations and ablation experiments on these datasets. Our system demonstrates superior performance compared to existing methods that scene reconstruction on urban scenes. In summary, our contributions are shown as follows:
@@
  • We propose a novel editable neural scene reconstruction method, ProSGNeRF, which decomposes dynamic, background, and far-field into a progressively neural scene graph with decoupled object transformations and scene representations and also supports object manipulation and scene roaming.
@@
  • We propose a progressive neural scene graph representation architecture for the large-scale scenes, We dynamically instantiate local neural scene graphs with temporal sliding windows, enabling scalability to handle arbitrary large-scale scenes and long sequences.
  • In each neural scene graph, the scene is decomposed into multiple dynamic objects and a background component. For background representation, we adopt a multiresolution hash encoding strategy to effectively capture information at different frequency levels, enabling accurate modeling of both geometric structure and color appearance at fine spatial scales.
@@
  • Due to the sparse observation of objects in the urban scene, we design a frequency auto-encoder network with vision foundation model to efficiently encode shape and appearance features for the objects of interest and perform frequency modulation for sparse view inputs. In addition, to better train our vision foundation model, we construct a curated dataset consisting of hundreds of thousands of 2D images and corresponding 3D spatial information from the KITTI and Waymo datasets.
  • We validate the effectiveness of our approach across multiple urban datasets, including KITTI and Virtual KITTI (vKITTI), Waymo datasets, which contain diverse scenarios with highly dynamic objects, as well as the large-scale Waymo dataset, which encompasses complex urban scenes with both static environments and dynamic objects such as vehicles and pedestrians.
Figure 2: The isometric view of the proposed method, ProSGNeRF. We employed a 2D segmentation network, SAM, to preprocess the training data and generate accurate masks for dynamic objects. We propose a progressive neural scene graph architecture that dynamically allocates local neural scene graph (box). The entire scene is decomposed into three parts: background, dynamic objects, and far-field. We design separate networks for background and objects and introduce a far-field loss for regularization. Nodes $l_{i}$ represent individual dynamic objects. $F_{bkg}$ models the static background scene and $F_{obj}$ models movable foreground objects in local object-centric coordinate frames.
@@ 2 Related Work
Recently, with the introduction and advancement of Neural Radiance Fields [NeRF], a broad array of subsequent research has been inspired, building upon the original approach. We review the related work in the areas of implicit scene representations, dynamics, and large-scale scene representation.
2.1 Implicit Scene Representation
@@
NeRF [NeRF] is the most impressive work in view synthesis and 3D reconstruction. It proposes implicit representations to model scenes as functions, encoding the scene using a multi-layer perceptron (MLP). NeRF++ [nerf++], Mip-nerf [mipnerf], and Mip-NeRF 360 [mipnerf360] further improve the representation ability for unbounded environments. Those methods use the non-linear scene parameterization method to model unbounded scenes. [srf] proposes a spherical radiance field (SRF) for efficient novel view synthesis in unbounded scenes. NeRF-W [nerfw] incorporate per-image embeddings to model appearance variation of the outdoor scenes. NeRF-Der++ [nerfdet] and Surface-SOS [surfacesos] combine implicit scene representation with object detection method for semantic and geometry consistency. To further accelerate the rendering process, 3DGS [3dgs] proposes to represent the entire scene using Gaussian primitives.
2.2 Dynamics
@@
There are some methods that focus on dynamic object reconstruction, such as Neural 3D Video Synthesis [n3dvs] and Space-time Neural radiance Fields [spacetime]. These approaches encode time information into scene representation. [nsff] and [dynerf] incorporate 2D optical flow with warping-based regularization losses to enforce multi-view consistency and dynamic object perception. Nerfies [nerfies] and Hypernerf [hypernerf] design a per-frame deformation field for dynamic humans. These methods are limited to single-object scenes, whereas our method scales beyond synthetic and indoor scenes to urban scenes.
2.3 Large-scale Scene Representation
@@
For large-scale scene representation, [bungeenerf] proposes Bungeenerf with a progressively growing model for city-level rendering. Block-NeRF [blocknerf] and Mega-NeRF [meganerf] decompose the scene into multiple local radiance fields for large-scale scenes. However, these methods can only deal with static environments. Dynamic objects are prevalent in real-world urban road scenes. Panoptic-NeRF [panopticnerf] and Panoptic Neural Fields [pnf] produce panoptic segmentation jointly with scene representations, achieving urban scene reconstruction.[incremental, plgslam] proposes incremental scene representation method for large-scale indoor environments.
@@
NSG [nsg], SUDS [suds], and MARS [mars] are the most relevant works to ours. NSG adopts neural scene graphs and dynamic object bounding boxes to decouple the scene but cannot handle ego-motion. Besides, NSG needs over 1TB of memory to represent a 30-second video. SUDS [suds] uses Foundation model [dino] to distill scene features and obtain semantic information and scene flow. Their methods can handle camera ego-motion and static objects but struggle with high-dynamic objects. With the proposal of 3D gaussian splating [3dgs] Some 3DGS-based methods, such as DrivingGaussian [drivinggaussian], StreetGaussian [streetgaussian] and [vpgsslam, pvg, mcnslam], also focus on urban scene reconstruction. However, the excessive memory consumption of 3DGS makes it unsuitable for large-scale scene reconstruction, and these approaches can only handle the reconstruction of 30–40 frames. In contrast, our approach effectively addresses large-scale camera ego-motion and dynamic objects and long video sequences, enabling urban scene editing useful for constructing simulators for autonomous driving.
@@ 3 Method
The pipeline of our system is shown in Fig. 2. We use a set of sequential RGB images $\{I_{i}\}_{i=1}^{M}$ with known pose $\{R|t\}$ as our input where $M$ indicates the number of images. We construct the progressive neural scene graph with the instance mask from SAM and 3D bounding box from machine-predicted. The scene is decomposed into multi-local scene graphs with three branches: background, dynamic objects, and far-field. In the dynamic objects branch, we design a frequency auto-encoder to learn a normalized, object-centric representation whose embedding describes and disentangles shape, appearance, and pose. The decoder translates the features into occupancy and color information for given 3D points and viewing directions in object space. We also design lidar points projection geometry consistency in large-scale urban scenes.
Figure 3: The isometric view of the proposed progressive neural scene graph. This scene representation uses a progressive scheme that dynamically allocates a local neural scene graph (box) based on the camera pose. Adjacent local scene graphs have overlapping regions to maintain global consistency. The leaf nodes $l$ are visualized as boxes with their local Cartesian coordinate axis. We also visualize the transformations and scaling between root and leaf coordinate frames using arrows with annotated transformation and scale matrices. The overall representation model is denoted as $F_{\theta}$.
@@ 3.1 Progressive Neural Scene Graph
To this end, we design a progressive neural scene graph to represent city-scale scenes with multi-dynamic objects. Instead of pre-dividing the scene with the camera poses, our method dynamically adjust local scene bounds during optimization, preventing the splitting of dynamic elements when applied to dynamic environments and improving the quality of reconstructing dynamic objects.
The scene graph $\mathrm{SG}$, illustrated in Fig. 3, is composed of cameras, static nodes and a set of dynamic nodes which represent the dynamic components of the scene, including the object’s appearance, shape, and classes. The entire scene is decomposed into several local scene representations:
$\{I_{i},P_{i}\}_{i=1}^{M}\mapsto\{\mathrm{SG}^{1}_{\theta_{1}},\mathrm{SG}^{2}_{\theta_{2}},\dots,\mathrm{SG}^{n}_{\theta_{n}}\}\mapsto\{c,\sigma\}$
@@
$P_{i}$ denotes the projection of the LiDAR point cloud at frame i. The root node of scene graph denoted as $\mathcal{W}$, $C$ is a leaf node representing the camera in the global coordinate system and its intrinsic $K\in\mathbb{R}^{3\times 3}$. $L$ denotes the leaf node of scene graph(e.g. objects and background). $F$ represents the neural networks for both static and dynamic nodes: $F=F_{\theta_{bckg}}\cup\left\{F_{\theta_{1}},F_{\theta_{2}},\dots,F_{\theta_{c}}\right\}$. $E$ are edges that represent affine transformations from $i$ to $i^{\prime}$:
$E$ $=\left\{e_{i,i^{\prime}}\in\left\{[],\boldsymbol{T}_{i\to i^{\prime}}\right\}\right\},$ (2)
$\text{ with }\quad\boldsymbol{T}_{i\to i^{\prime}}$ $=\left[\begin{array}[]{cc}\boldsymbol{R}_{i\to i^{\prime}}&\boldsymbol{t}_{i\to i^{\prime}}\\ \mathbf{0}&1\end{array}\right],$
$\boldsymbol{R}_{i\to i^{\prime}}$ $\in\mathrm{SO(3)},\quad\boldsymbol{t}_{i\to i^{\prime}}\in\mathrm{SE(3)}{.}$
For all edges originating from the root node $\mathcal{W}$, we assign transformations between the global world space and the local objects with $T$. We compute the scaling factor $S_{i}$, assigned at edge $e_{i}$ between an object node and its shared representation model $F$.
We present our progressive scheme in Algorithm 1. The algorithm describes a method for constructing a Progressive Neural Scene Graph that dynamically allocates a new local scene graph trained with frames within a temporal window, allowing us to scale up the representation to an arbitrarily large scene. Whenever the estimated camera pose trajectory reaches the bound of the current scene graph. we initialize the local representation with a local subset of frames, and then we progressively introduce subsequent frames to the optimization. For each frame, it optimizes the local neural scene graph and progresses to the next frame until a predefined condition is met. Each subset contains some overlapping images, essential for achieving consistent reconstructions in the global scene. When the limit for a local neural scene graph is reached, it initializes a new one, centered around the last pose, and continues the process until all frames are processed, contributing to the complete scene reconstruction. We further use inverse distance weight (IDW) fusion for all overlapping scene representations. Whenever the camera pose leaves the bound of the current scene representation, we stop optimizing previous ones (freeze the network parameters). At this point, we can reduce memory requirements and clear any supervising frames that are not needed anymore.
Algorithm 1 Our Progressive Neural Scene Graph.
@@
[…]
6:while $q<P$ do
7:  while $\{R,t\}_{j}<Bound_{j}$ do
8:   $q\leftarrow q+1$;
9:   $\{R,T\}_{q}\leftarrow\{R,T\}_{q-1}$; $\ast$ Append a pose at the end of the trajectory $\ast/$
10:   Optimize($\theta_{j}$); $\ast$ Optimize local neural scene graph $\ast/$
11:  end
12:  if $q<P$ then
[…]
17:  end if
18:  Repeat until all frames are registered.
19:end
Figure 4: Ray-Box Intersection. We use AABB ray-box sampling strategy. The boxes are now defined as an axis-aligned bounding box (AABB) with a minimum bound [-1,-1,-1] and a maximum bound [1, 1,1].
Table 1: Background Model Architecture.
Background Model First Stage $F_{bckg,1}$
Layer Name Input Size
Input Local Position + Fourier 3 + 60
layer1 Fully Connected ReLU 256
layer2 Fully Connected ReLU 256
layer3 Fully Connected ReLU 256
layer4 Fully Connected ReLU + Skip 256
layer5 Fully Connected ReLU 256
layer6 Fully Connected ReLU 256
layer7 Fully Connected ReLU 256
PrevOut, Density Fully Connected ReLU 256
Background Model Second Stage $F_{bckg,2}$
Layer Name Input Size
Input Direction + Fourier + PrevOut 3 + 24 + 256
second1 Fully Connected ReLU 256
second2 Fully Connected ReLU 256
second3 Fully Connected ReLU 256
RGB Fully Connected ReLU 256